# Lighter, Better, Faster Multi-Source Domain Adaptation with Gaussian Mixture Models and Optimal Transport

### v2

@@ 3.3 Mixture Wasserstein Barycenters
In this section, we detail our first methodological contribution: a new algorithm for computing barycenters of GMMs under the $\mathcal{MW}_{2}$ and $\mathcal{SMW}_{2}$ metrics. As such, we adapt the definition of [21],
@@
Given $C\geq 1$ GMMs $\mathcal{P}=\{P_{c}\}_{c=1}^{C}$ and a vector of barycentric coordinates $\lambda\in\Delta_{C}$, the $\mathcal{SMW}_{2}$ barycenter is given by,
$P^{\star}=\mathcal{B}(\lambda,\mathcal{P})=\underset{B\in\text{GMM}_{d}(\infty)}{\text{argmin}}\,\biggr{\{}\mathcal{L}(B)=\sum_{c=1}^{C}\lambda_{c}\mathcal{SMW}_{2}(B,P_{c})^{2}\biggr{\}}.$ (14)
When the GMMs in $\mathcal{P}$ are unlabeled, one may define, by analogy, a barycenter under the $\mathcal{MW}_{2}$. Henceforth we describe an algorithm for labeled GMMs, but its extension for unlabeled GMMs is straightforward.
Inspired by previous results in empirical Wasserstein barycenters [22, 12], we propose a novel strategy for computing $\mathcal{B}(\lambda,\mathcal{P})$. Our idea consists on initializing the parameters $\{(\mathbf{m}_{i}^{(B)},\mathbf{s}_{i}^{(B)},\mathbf{y}_{i}^{(B)})\}_{i=1}^{K_{B}}$ at random. We then calculate $C$ GMM-OT transport plans $\{\omega^{(c)}\}_{c=1}^{C}$ between $P_{c}$ and $B$. Next, we use equations 10 and 13 to update the parameters of the barycenter $B$. We refer readers to our supplementary materials for further details. This strategy effectively minimizes $\mathcal{L}$ in eq. 14. We show our strategy in Algorithm 1.
1 function mixture_wasserstein_barycenter($\{\mathbf{M}^{(P_{c})}\}_{c=1}^{C}$, $\{\mathbf{S}^{(P_{c})}\}_{c=1}^{C}$, $\{\mathbf{Y}^{(P_{c})}\}_{c=1}^{C}$, $\{\lambda_{c}\}_{c=1}^{C}$, $\tau$, $N_{it}$)
2       $\mathbf{m}_{i}^{(B)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$, $\mathbf{s}_{i}^{(B)}:=1$ and $\mathbf{y}_{i}^{(B)}:=\nicefrac{{1}}{{n_{c}}}$
@@ 5                   $\omega^{(c,it)}=\text{GMMOT}(B,P_{c})$
7             $L_{it}=\sum_{c=1}^{C}\lambda_{c}\sum_{i=1}^{K_{B}}\sum_{j=1}^{K_{P}}\omega^{(c,it)}_{ij}\biggr{(}(\mathcal{W}_{2}(B_{i},P_{c,j})^{2}+\lVert\mathbf{y}^{(B)}_{i}-\mathbf{y}^{(P_{c})}_{j}\rVert_{2}^{2})\biggr{)}$
@@
8             $\mathbf{m}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{k}T_{\omega^{(k,it)}}(\mathbf{m}^{(B)}_{i})$
9             $\mathbf{s}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{k}T_{\omega^{(k,it)}}(\mathbf{s}^{(B)}_{i})$
10             $\mathbf{y}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{k}T_{\omega^{(k,it)}}(\mathbf{y}^{(B)}_{i})$
@@
12      return $\mathbf{M}^{(B)}$, $\mathbf{S}^{(B)}$, $\mathbf{Y}^{(B)}$
@@ 3.4 Multi-Source Domain Adaptation through GMM-OT
In this section, we detail two contributions for MSDA based on GMM-OT: GMM-WBT and GMM-DaDiL. In both cases, we suppose access to $N$ labeled source measures $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N}$ and an unlabeled target measure $Q_{T}$. Contrary to the empirical versions of these algorithms [8, 9, 12], we assume that an axis-aligned GMM has been learned for each domain, including the target.
@@
With a labeled GMM, $\{\hat{\mathbf{m}}_{i}^{(Q_{T})},\hat{\mathbf{s}}_{i}^{(Q_{T})},\mathbf{y}_{i}^{(B)}\}_{i=1}^{K_{B}}$, on the target domain, we can learn a classifier on the target domain as explained in section 3.2.
1 function gmm_dadil($\{\mathbf{M}^{(Q_{S_{\ell}})}\}_{\ell=1}^{N}$, $\{\mathbf{S}^{(Q_{S_{\ell}})}\}_{\ell=1}^{N}$, $\{\mathbf{Y}^{(Q_{S_{\ell}})}\}_{\ell=1}^{N}$, $N_{it}$, $\eta$)
@@
4             $L=0$
@@ 8             $L\leftarrow L+\mathcal{MW}_{2}(Q_{T},\mathcal{B}(\lambda_{T},\mathcal{P}))^{2}$
9             $\mathbf{x}_{j}^{(P_{k})}\leftarrow\mathbf{x}_{j}^{(P_{k})}-\eta\nicefrac{{\partial L}}{{\partial\mathbf{x}^{(P_{k})}_{j}}}$
10             $\mathbf{p}_{j}^{(P_{k})}\leftarrow\mathbf{u}_{j}^{(P_{k})}-\eta\nicefrac{{\partial L}}{{\partial\mathbf{u}^{(P_{k})}_{j}}}$
@@
12             $\lambda_{\ell}\leftarrow\text{proj}_{\Delta_{K}}(\lambda_{\ell}-\eta\nicefrac{{\partial L}}{{\partial\lambda_{\ell}}})$
13
14      return $\mathcal{P},\Lambda$
15
@@
GMM-DaDiL. Our second algorithm consists of a parametric version for the DaDiL algorithm of [12]. The idea is to replace the atoms in $\mathcal{P}=\{\hat{P}_{c}\}_{c=1}^{C}$ by GMMs parametrized through $\Theta_{P}=\{(\mathbf{M}^{(P_{c})},\mathbf{S}^{(P_{c})},\mathbf{Y}^{(P_{c})})\}_{c=1}^{C}$. Learning a dictionary is thus equivalent to estimating these parameters, that is,
$\mathcal{L}(\Lambda,\Theta_{P})$ $=\underset{\Lambda,\mathcal{P}}{\text{argmin}}\,\mathcal{MW}_{2}(Q_{T},\mathcal{B}(\lambda_{T};\mathcal{P}))^{2}+\sum_{\ell=1}^{N}\mathcal{SMW}_{2}(Q_{\ell},\mathcal{B}(\lambda_{\ell};\mathcal{P}))^{2}.$ (16)
An important feature of this strategy is that the barycenters $\mathcal{B}$ inherit the labels from the atoms $\mathcal{P}$. As a result, we can estimate a labeled GMM for $Q_{T}$ in one step through $\mathcal{B}(\lambda_{T};\mathcal{P})$.
@@
Once the dictionary $(\Lambda,\mathcal{P})$ is learned, we are able to reconstruct the domains in MSDA via the barycenter $\mathcal{B}(\lambda;\mathcal{P})$. We are especially interested in the target reconstruction $\lambda_{T}$, i.e., $\mathcal{B}(\lambda_{T};\mathcal{P})$. This barycenter is a labeled GMM (as we show in figure 5b). As a result, we can obtain labeled samples from this GMM, then use them to train a classifier that works on the target domain.
Faster Dataset Dictionary Learning. The computational complexity of an optimization step of algorithm 2 corresponds to $\mathcal{O}(N\times N_{it}\times C\times K^{3}\log K)$, i.e., we calculate $N$ barycenters of $C$ atoms. One should compare this complexity with that of DaDiL, i.e., $\mathcal{O}(N\times N_{it}\times M\times C\times n_{b}^{3}\log n_{b})$, where $M=\lceil\nicefrac{{n}}{{n_{b}}}\rceil$ is the number of mini-batches sampled at each iteration. In our experiments in section 4, we show that we achieve state-of-the-art performance with $K$ on the same order of magnitude as $n_{b}$ (e.g., a few hundred Gaussian components). As a result, we achieve a speed-up on the order of $M$ while solving an exact OT problem.
@@ 4 Experiments
In this section, we present our experiments. First, we present a toy example, where the source and target measures were generated by transforming an initial measure. As such, this toy example illustrates our goal to model distributional shift. Second, we experiment with benchmarks in MSDA.
Figure 3: Data used in the toy experiment. Each of these datasets was generated by applying an affine transformation to an initial dataset.
Toy Example. In figure 3 we show the generated datasets. Starting with GMM-WBT, figure 4(a) shows the learned GMMs for each dataset, in which the target GMM is not labeled. The barycenter of $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N}$ is shown in figure 4(b) (bottom-left). This barycenter is labeled. As a result, we may transfer its parameters to the target domain through GMM-OT (upper part), which leads to a labeled GMM in the target domain.
(a) GMMs learned on each domain.
(b) GMM-WBT.
Figure 4: Summary of GMM-WBT, in which a labeled GMM is determined on the target domain by transporting the parameters of the barycenter of sources.
@@
Figure 5: Optimization of GMM-DaDiL and reconstruction of domain GMMs.
Experimental Setting. We compare our method to prior art. We focus on OT-based methods, such as WJDOT [10], WBT [8, 9] and DaDiL [12]. For completeness, we include recent strategies on deep MSDA that update the encoder network during the adaptation process, rather than using pre-extracted features. These are M3SDA [23], LtC-MSDA [24], KD3A [25] and Co-MDA [26]. We establish our comparison on 4 benchmarks, divided between visual domain adaptation (Office31 [27], Office-Home [28]) and cross-domain fault diagnosis (TEP [29] and CWRU). See table 1 for further details.
@@
Benchmarking Results. First, OT-based methods generally outperform other methods in MSDA. Overall, these methods solve a simpler task compared to deep DA methods, as they do not need to update the encoder network during adaptation. Second, the GMM-OT framework generally improves over using empirical OT. For instance, in the CWRU benchmark, GMM-WBT largely outperforms WBT [8, 9]. Furthermore, GMM-DaDiL outperforms its empirical counterpart on all benchmarks, as well as GMM-WBT. This point further showcases the power of dictionary learning in MSDA.
@@
Figure 6: In (a), we analyse the performance of interpolations $\mathcal{B}((\lambda_{0},1-\lambda_{0});\mathcal{Q}_{S})$, $\mathcal{Q}_{S}=\{Q_{S_{1}},Q_{S_{2}}\}$ and $\mathcal{B}((\lambda_{0},1-\lambda_{0});\mathcal{P})$ with learned $\mathcal{P}=\{P_{1},P_{2}\}$ for GMM-WBT and GMM-DaDiL. In (b), we analyse the efficiency of barycenter-based methods under an increasing number of GMM components (number of samples for DaDiL and WBT). GMM-DaDiL has state-of-the-art performance even for the extreme case where $K=65$.
Better MSDA. We design an experiment to further showcase the contribution of dictionary learning for MSDA. We use the adaptation $(D,W)\rightarrow A$ in the Office-31 benchmark as the basis of our experiment. Note that GMM-DaDiL reconstructs the target domain via a barycenter $\mathcal{B}(\lambda_{T};\mathcal{P})$, where $\lambda_{T}$ and $\mathcal{P}$ are learned parameters. We thus ablate the learning of $\lambda_{T}$ and $\mathcal{P}$, i.e., we compare it to $\lambda_{T}=(\lambda_{0},1-\lambda_{0})$, $\lambda_{0}\in[0,1]$ and $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N_{S}}$. This generates a series of measures parametrized by $\lambda_{0}$. To further match $\mathcal{B}(\lambda_{T};\mathcal{Q}_{S})$ with $Q_{T}$, we transport it to $Q_{T}$ through eq. 15. Note that this corresponds to performing GMM-WBT with a barycenter calculated with $\lambda_{T}$. The overall experiment is shown in figure 6. While the performance of GMM-WBT remains approximately stable, that of GMM-DaDiL grows as we move closer to $\lambda_{T}^{\star}$ learned by dictionary learning (blue star). Overall, the interpolation space generated by atoms better captures the distributional shift occurring on the target domain.
@@
Lighter MSDA. We further investigate the adaptation efficiency of our methods with respect the number of GMM components $n$. We use the adaptation task $(Cl,Pr,Rw)\rightarrow Ar$ from Office-Home to showcase our analysis. Note that this parameter is analogous to the number of samples in DaDiL or WBT. We thus show a comparison in figure 6 b. From this figure, we see that GMM-DaDiL surpasses all other methods over the entire range $n\in\{65,130,\cdots,910\}$, which showcases the advantage of using GMMs and dictionary learning. Especially, its empirical counterpart, DaDiL, needs a large number of samples in order to accurately represent probability measures.
Curiously, the performance of GMM-WBT and WBT are quite similar. Indeed, recent studies [31] show that Wasserstein barycenters are effective in compressing probability measures with respect the number of their samples. As a result, in this adaptation task, the GMM version of WBT has similar performance to the empirical version.
@@
Visualization. In this experiment, we use the $(B,C)\rightarrow A$ adaptation task of CWRU. We are interested in visualizing the evolution of atoms and reconstructions with respect DaDiL and GMM-DaDiL iterations. We visualize this evolution through UMAP [32], i.e., we concatenate the data from DaDiL’s atoms, i.e., $\mathbf{x}_{i}^{(P_{c,it})}$, so that these are jointly embedded into $\mathbb{R}^{2}$. For GMM-DaDiL, we concatenate the mean parameters, i.e., $\mathbf{m}_{i}^{(P_{c,it})}$. We summarize our results in figure 7. Overall, as shown in figure 7 (a – d), GMM-DaDiL optimization is more stable than that of DaDiL, especially since we do not use mini-batches. This remark is also evidenced in the reconstructions in figure 7 (e – f).

### v3

@@ 3.3 Mixture Wasserstein Barycenters
In this section, we detail a new algorithm for computing barycenters of GMMs under the $\mathcal{MW}_{2}$ and $\mathcal{SMW}_{2}$ metrics. As such, we adapt the definition of [21],
@@
Given $C\geq 1$ GMMs $\mathcal{P}=\{P_{c}\}_{c=1}^{C}$, $K_{B}\geq 1$, and a vector of barycentric coordinates $\lambda\in\Delta_{C}$, the $\mathcal{SMW}_{2}$ barycenter is given by,
$B^{\star}=\mathcal{B}(\lambda,\mathcal{P})=\underset{B\in\text{GMM}_{d}(K_{B})}{\text{argmin}}\,\biggr{\{}\mathcal{L}(B)=\sum_{c=1}^{C}\lambda_{c}\mathcal{SMW}_{2}(B,P_{c})^{2}\biggr{\}}.$ (14)
When the GMMs in $\mathcal{P}$ are unlabeled, one may define, by analogy, a barycenter under the $\mathcal{MW}_{2}$. Henceforth we describe an algorithm for labeled GMMs, but its extension for unlabeled GMMs is straightforward. Inspired by previous results in empirical Wasserstein barycenters [22, 12], we propose a novel strategy for computing $\mathcal{B}(\lambda,\mathcal{P})$. Our method relies on the analysis of $\theta_{B}=\{(\mathbf{m}_{i}^{(B)},\mathbf{s}_{i}^{(B)},\mathbf{v}_{i}^{(B)})\}_{i=1}^{K_{B}}\mapsto\sum_{c=1}^{C}\lambda_{c}\mathcal{SMW}_{2}(B,P_{c})^{2}$. First, for a fixed $\theta_{B}$, we find $\omega_{1}^{\star},\cdots,\omega_{C}^{\star}$ transport plans. Then, for fixed transport plans, we solve,
$\underset{\theta_{B}}{\text{argmin}}\,\mathcal{L}(\theta_{B})$ $=\sum_{c=1}^{C}\lambda_{c}\sum_{i=1}^{K_{B}}\sum_{j=1}^{K_{P}}\omega_{c,i,j}^{\star}C_{c,i,j},$
where $\text{where }C_{c,i,j}$ $=\lVert\mathbf{m}_{i}^{(B)}-\mathbf{m}_{j}^{(P_{c})}\rVert_{2}^{2}+\lVert\mathbf{s}_{i}^{(B)}-\mathbf{s}_{j}^{(P_{c})}\rVert_{2}^{2}+\beta\lVert\mathbf{v}_{i}^{(B)}-\mathbf{v}_{j}^{(P_{c})}\rVert_{2}^{2}$
which can be optimized by taking derivatives with respect $\mathbf{m}^{(B)}_{i}$, $\mathbf{s}^{(B)}_{i}$ and $\mathbf{v}^{(B)}_{i}$. For instance, taking the derivative of $\mathcal{L}(\theta_{B})$ with respect $\mathbf{m}_{i}^{(B)}$,
setting the derivative to 00, one has, $\mathbf{m}_{i}^{(B)}=\sum_{c=1}^{C}\lambda_{c}T_{\omega_{c}^{\star}}(\mathbf{m}_{i}^{(B)}).$ Similar results can be acquired for $\mathbf{s}_{i}^{(B)}$ and $\mathbf{v}_{i}^{(B)}$ by taking the appropriate derivatives. Our strategy is shown in Algorithm  1.
1 function smw_barycenter($\{(\mathbf{M}^{(P_{c})},\mathbf{S}^{(P_{c})},\mathbf{V}^{(P_{c})})\}_{c=1}^{C}$, $\tau$, $N_{it}$)
2       $\mathbf{m}_{i}^{(B)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$, $\mathbf{s}_{i}^{(B)}=1$ and $\mathbf{y}_{i}^{(B)}=\nicefrac{{\mathds{1}_{n_{c}}}}{{n_{c}}}$
@@ 5                   $\omega^{(c,it)}=\text{GMMOT}(B,P_{c})$
7             $L_{it}=\sum_{c=1}^{C}\lambda_{c}\sum_{i=1}^{K_{B}}\sum_{j=1}^{K_{P}}\omega^{(c,it)}_{ij}\biggr{(}(\mathcal{W}_{2}(B_{i},P_{c,j})^{2}+\beta\lVert\mathbf{v}^{(B)}_{i}-\mathbf{v}^{(P_{c})}_{j}\rVert_{2}^{2})\biggr{)}$
@@
8             $\mathbf{m}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{c}T_{\omega^{(c,it)}}(\mathbf{m}^{(B)}_{i})$
9             $\mathbf{s}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{c}T_{\omega^{(c,it)}}(\mathbf{s}^{(B)}_{i})$
10             $\mathbf{v}^{(B)}_{i}=\sum_{c=1}^{C}\lambda_{c}T_{\omega^{(c,it)}}(\mathbf{v}^{(B)}_{i})$
@@
12      return $\mathbf{M}^{(B)}$, $\mathbf{S}^{(B)}$, $\mathbf{V}^{(B)}$
@@ 3.4 Multi-Source Domain Adaptation through GMM-OT
In this section, we detail two contributions for MSDA based on GMM-OT: GMM-WBT and GMM-DaDiL. In both cases, we suppose access to $N$ labeled source GMMs $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N}$ and an unlabeled target GMM $Q_{T}$. Contrary to the empirical versions of these algorithms [8, 9, 12], we assume that an axis-aligned GMM has been learned for each domain, including the target.
@@
With a labeled GMM, $\{\hat{\mathbf{m}}_{i}^{(Q_{T})},\hat{\mathbf{s}}_{i}^{(Q_{T})},\mathbf{v}_{i}^{(B)}\}_{i=1}^{K_{B}}$, on the target domain, we can learn a classifier on the target domain as explained in section 3.2.
1 function gmm_dadil($\{(\mathbf{M}^{(Q_{S_{\ell}})},\mathbf{S}^{(Q_{S_{\ell}})},\mathbf{V}^{(Q_{S_{\ell}})})\}_{\ell=1}^{N}$, $\{(\mathbf{M}^{(Q_{T})},\mathbf{S}^{(Q_{T})})\}$, $N_{it}$, $\eta$)
@@
4             $L\leftarrow 0$
             // Change of variables
5             $\mathbf{v}_{i}^{(P_{k})}\leftarrow\text{softmax}(\mathbf{u}_{i}^{(P_{k})})$
@@ 9             $L\leftarrow L+\mathcal{MW}_{2}(Q_{T},\mathcal{B}(\lambda_{T},\mathcal{P}))^{2}$
10             $\mathbf{m}_{j}^{(P_{k})}\leftarrow\mathbf{m}_{j}^{(P_{k})}-\eta\nicefrac{{\partial L}}{{\partial\mathbf{m}^{(P_{k})}_{j}}}$
11             $\mathbf{u}_{j}^{(P_{k})}\leftarrow\mathbf{u}_{j}^{(P_{k})}-\eta\nicefrac{{\partial L}}{{\partial\mathbf{u}^{(P_{k})}_{j}}}$
@@
13             $\lambda_{\ell}\leftarrow\text{proj}_{\Delta_{C}}(\lambda_{\ell}-\eta\nicefrac{{\partial L}}{{\partial\lambda_{\ell}}})$
14
15      return $\Lambda,\mathcal{P}$
16
@@
GMM-DaDiL. Our second algorithm consists of a parametric version for the DaDiL algorithm of [12]. The idea is to replace the atoms in $\mathcal{P}=\{\hat{P}_{c}\}_{c=1}^{C}$ by GMMs parametrized through $\Theta_{P}=\{(\mathbf{M}^{(P_{c})},\mathbf{S}^{(P_{c})},\mathbf{V}^{(P_{c})})\}_{c=1}^{C}$. Learning a dictionary is thus equivalent to estimating these parameters, that is,
$(\Lambda^{\star},\Theta_{P}^{\star})$ $=\underset{\Lambda,\Theta_{P}}{\text{argmin}}\,\mathcal{MW}_{2}(Q_{T},\mathcal{B}(\lambda_{T},\mathcal{P}))^{2}+\sum_{\ell=1}^{N}\mathcal{SMW}_{2}(Q_{\ell},\mathcal{B}(\lambda_{\ell};\mathcal{P}))^{2}.$ (16)
@@
Once the dictionary $(\Lambda,\mathcal{P})$ is learned, we are able to reconstruct the domains in MSDA via the barycenter $\mathcal{B}(\lambda;\mathcal{P})$. We are especially interested in the target reconstruction $\lambda_{T}$, i.e., $\mathcal{B}(\lambda_{T},\mathcal{P})$. This barycenter is a labeled GMM (as we show in figure 4b). As a result, we can obtain labeled samples from this GMM, then use them to train a classifier that works on the target domain.
The computational complexity of an optimization step of algorithm 2 corresponds to $\mathcal{O}(N\times N_{it}\times C\times K^{3}\log K)$, i.e., we calculate $N$ barycenters of $C$ atoms. One should compare this complexity with that of DaDiL, i.e., $\mathcal{O}(N\times N_{it}\times M\times C\times n_{b}^{3}\log n_{b})$, where $M=\lceil\nicefrac{{n}}{{n_{b}}}\rceil$ is the number of mini-batches sampled at each iteration. In our experiments in section 4, we show that we achieve state-of-the-art performance with $K$ on the same order of magnitude as $n_{b}$ (e.g., a few hundred Gaussian components). As a result, we achieve a speed-up on the order of $M$ while solving an exact OT problem (see Figure 5 (c) below).
@@ 4 Experiments
4.1 Toy Example
In this section, we explore GMM-WBT and GMM-DaDiL in the context of a toy example. We generate 4 datasets over $\mathbb{R}^{d}$, by gradually shifting and deforming initial measure through an affine mapping. In figure 3 (a) we show the generated datasets. Starting with GMM-WBT, figure 3 (b) shows the learned GMMs for each dataset, in which the target GMM is not labeled. The barycenter of $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N}$ is shown in figure 3(c) (bottom-left). This barycenter is labeled. As a result, we may transfer its parameters to the target domain through GMM-OT (upper part), which leads to a labeled GMM in the target domain.
(a) Data.
(b) GMMs.
(c) GMM-WBT
Figure 3: Data and GMMs used in the toy experiment. In (a) Each of these datasets was generated by applying an affine transformation to an initial dataset. In (b), we show an axis-aligned GMM fitted to the data via EM. In (c), we show a summary of GMM-WBT, where show the OT plan between components (upper part) between $B$ (left) and $Q_{T}$ (right). The resulting labeled GMM is shown in the lower right part of (c).
@@ 4.1 Toy Example
Figure 4: Optimization and reconstruction summaries using GMM-DaDiL. In (a), we show the evolution of loss, negative log-likelihood and barycentric coordinates (i.e., $\lambda_{\ell}$) over the course of optimization. In (b), we show the reconstructed GMMs (i.e., $\mathcal{B}(\lambda_{\ell},\mathcal{P})$) when the algorithm converges.
4.2 Multi-Source Domain Adaptation
We compare our method to prior art. We focus on OT-based methods, such as WJDOT [10], WBT [8, 9] and DaDiL [12]. For completeness, we include recent strategies on deep MSDA that update the encoder network during the adaptation process, rather than using pre-extracted features. These are M3SDA [23], LtC-MSDA [24], KD3A [25] and Co-MDA [26]. We establish our comparison on 4 benchmarks, divided between visual domain adaptation (Office31 [27], Office-Home [28]) and cross-domain fault diagnosis (TEP [29] and CWRU). See table 1 for further details.
@@ 4.2 Multi-Source Domain Adaptation
First, OT-based methods generally outperform other methods in MSDA. Overall, shallow DA methods solve a simpler task compared to deep DA methods, as they do not need to update the encoder network during adaptation. Second, the GMM-OT framework generally improves over using empirical OT. For instance, in the CWRU benchmark, GMM-WBT largely outperforms WBT [8, 9]. Furthermore, GMM-DaDiL outperforms its empirical counterpart on all benchmarks, as well as GMM-WBT. This point further illustrates the power of dictionary learning in MSDA. Note that, in table 2(d) (d), GMM-DaDiL manages to have the best average adaptation performance across domains without actually being the best on any single domain. As a consequence, GMM-DaDiL enjoys better stability, with respect distribution shift, than previous methods.
4.3 Lighter, Better, Faster Domain Adaptation
(a) Lighter.
(b) Better.
(c) Faster.
@@ 4.3 Lighter, Better, Faster Domain Adaptation
Figure 5: Lighter, Better, Faster. In (a), we analyse the performance of interpolations $\mathcal{B}((\lambda_{0},1-\lambda_{0});\mathcal{Q}_{S})$, $\mathcal{Q}_{S}=\{Q_{S_{1}},Q_{S_{2}}\}$ and $\mathcal{B}((\lambda_{0},1-\lambda_{0});\mathcal{P})$ with learned $\mathcal{P}=\{P_{1},P_{2}\}$ for GMM-WBT and GMM-DaDiL. In (b), we analyse the efficiency of barycenter-based methods under an increasing number of GMM components (number of samples for DaDiL and WBT). GMM-DaDiL has state-of-the-art performance even for the extreme case where $K=65$. In (c), we compare the running time of GMM-DaDiL with that of DaDiL, as a function of number of components $K$ and batch size $n_{b}$, respectively. This figure illustrates the speedup of GMM-DaDiL as the number of samples in DaDiL (and hence, $M=\lceil\nicefrac{{n}}{{n_{b}}}\rceil$) increases. Circles represent the average over $5$ independent runs, while the error bars show $2$ times the standard deviation.
@@
Our first experiment illustrates why GMM-DaDiL is lighter than previous barycenter-based algorithms, such as DaDiL. In this context, a lighter model needs less parameters to achieve a certain domain adaptation performance. We rank GMM-OT models by the number of components $K$, and empirical models by the number of samples $n$ in their support. Note that these parameters regulate the complexity of these algorithms. We use the adaptation task $(Cl,Pr$,$Rw)\rightarrow Ar$ from Office-Home for our analysis. We show a comparison in figure 5 (a). From this figure, we see that GMM-DaDiL surpasses all other methods over the entire range $K\in\{65,130,\cdots,910\}$. Especially, its empirical counterpart, DaDiL, needs a large number of samples for accurately represent probability measures. Curiously, the performance of GMM-WBT and WBT are quite similar. Indeed, recent studies [32] show that Wasserstein barycenters are effective in compressing probability measures with respect the number of their samples. As a result, in this adaptation task, the GMM version of WBT has similar performance to the empirical version.
Our second experiment illustrates why GMM-Optimal Transport Domain Adaptation (OTDA) provides a better framework for MSDA. We use the adaptation $(D,W)\rightarrow A$ in the Office-31 benchmark as the basis of our experiment. Note that GMM-DaDiL reconstructs the target domain via a barycenter $\mathcal{B}(\lambda_{T},\mathcal{P})$, where $\lambda_{T}$ and $\mathcal{P}$ are learned parameters. We thus ablate the learning of $\lambda_{T}$ and $\mathcal{P}$, i.e., we compare it to $\lambda_{T}=(\lambda_{0},1-\lambda_{0})$, $\lambda_{0}\in[0,1]$ and $\mathcal{Q}_{S}=\{Q_{S_{\ell}}\}_{\ell=1}^{N_{S}}$. This generates a series of measures parametrized by $\lambda_{0}$. To further match $\mathcal{B}(\lambda_{T},\mathcal{Q}_{S})$ with $Q_{T}$, we transport it to $Q_{T}$ through eq. 15. Note that this corresponds to performing GMM-WBT with a barycenter calculated with $\lambda_{T}$. The overall experiment is shown in figure 5 (b). While the performance of GMM-WBT remains approximately stable, that of GMM-DaDiL grows as we move closer to $\lambda_{T}^{\star}$ learned by dictionary learning (blue star). Overall, the interpolation space generated by atoms better captures the distributional shift occurring on the target domain.
Our third experiment shows that GMM-DaDiL is faster than DaDiL. We plot the running time of these methods on the Office 31 benchmark, for the $(D,W)\rightarrow A$ adaptation task. The variables that influence the complexity of GMM and empirical DaDiL are the number of components $K$ and the batch size $n_{b}$, respectively. For GMM-DaDiL, we simply measure its running time for $5$ independent runs of the algorithm (blue curve) for each $K\in\{31,62,\cdots,217\}$. For DaDiL, we set $n_{b}\in\{31,62,\cdots,217\}$, and set $n=M\times n_{b}$, where $M$ is the number of mini-batches. We measure the performance over $5$ independent runs as well. Other than these parameters, we fix $N_{iter}=50$ and $C=3$. As shown in Figure 5 (c), the running time of GMM-DaDiL and DaDiL are essentially equivalent for $M=1$. For $M>1$, we have a speedup that is proportional to $M$.
In our fourth experiment, we use the $(B,C)\rightarrow A$ adaptation task of CWRU. We are interested in visualizing the evolution of atoms and reconstructions with respect DaDiL and GMM-DaDiL iterations. We visualize this evolution through UMAP [33], i.e., we concatenate the data from DaDiL’s atoms, i.e., $\mathbf{x}_{i}^{(P_{c,it})}$, so that these are jointly embedded into $\mathbb{R}^{2}$. For GMM-DaDiL, we concatenate the mean parameters, i.e., $\mathbf{m}_{i}^{(P_{c,it})}$. We summarize our results in figure 6. Overall, as shown in figure 6 (a – d), GMM-DaDiL optimization is more stable than that of DaDiL, especially since we do not use mini-batches. This remark is also evidenced in the reconstructions in figure 6 (e – f).